Solve by the elimination method
2x-6y=2
2x-6y=3
Solve by the subsititution method.
4m+4=30
m-4n=33

Answer by Alan3354(69443)

Solve by the elimination method
2x-6y=2
2x-6y=3
This is inconsistent, 2x-6y can't equal both 2 and 3.
--------------
Solve by the subsititution method.
4m+4=30 -->2m+2 = 15
2m = 13
m = 13/2
------------
4n = 33-m = 33 - 13/2 = 53/2
n = 53/8